# Zero modes and index theorems for non-Hermitian Dirac fermions

### v1

@@ Abstract
Dirac fermions, subject to external magnetic fields and in the presence of mass orders that assume topologically nontrivial spatial textures such as domain-wall and vortices, for example, bind robust mid-gap states at zero-energy, the number of which is governed by the Aharonov-Casher and Jackiw-Rebbi or Jackiw-Rossi index theorems, respectively. Here I extend the jurisdiction of these prominent index theorems to Lorentz invariant non-Hermitian (NH) Dirac operators, constructed by augmenting the celebrated Dirac Hamiltonian by a masslike anti-Hermitian operator that also scales linearly with momentum. The resulting NH Dirac operator manifests real eigenvalues over an extended NH parameter regime, characterized by a real effective Fermi velocity for NH Dirac fermions. From the explicit solutions of the zero-energy bound states, I show that in the presence of external magnetic fields of arbitrary shape such modes always exist when the system encloses a finite number of magnetic flux quanta, while in the presence of spatially non-trivial textures of the mass orders localized zero-energy modes can only be found in the spectrum when the effective Fermi velocity for NH Dirac fermions is real. These findings pave a concrete route to realize nucleation of competing orders from the topologically robust zero-energy manifold in NH or open Dirac systems. Possible experimental setups to test these predictions are discussed.
@@ I Introduction
Zero energy in the Dirac theory enjoys a special status as the associated eigenvalue spectrum extends equally to positive and negative values [1, 2, 3]. Therefore, any bound state that resides precisely at zero-energy often (if not always) is robust due to the intrinsic particle-hole symmetry in the system. Such a situation occurs at least in two specific cases. (a) When massless Dirac fermions, confined to an Euclidean plane, are subject to external uniform or inhomogeneous magnetic fields, the number of zero-energy bound states is exactly equal to the total magnetic field quanta enclosed by the system, guaranteed by the Aharonov-Casher index theorem [4, 5, 6]. (b) Bound states at zero-energy appear in the spectrum of Dirac fermions when they foster mass orders, captured by Hermitian operators fully anticommuting with the Dirac Hamiltonian, that assume spatially non-trivial topological textures. For example, in one spatial dimension, when the Dirac mass takes the shape of a domain-wall a zero-energy mode gets pinned where the mass order changes sign, a result known as the Jackiw-Rebbi index theorem [7, 8, 9]. In two spatial dimensions, when a composite Dirac mass, described by two mutually anticommuting mass matrices with the requisite U(1) symmetry between them, takes the texture of a vortex, $n$ number of zero-energy bound states get pinned near the core of such topological defects, where $n$ is the vorticity of the mass texture. Such a one-to-one correspondence between the vorticity of the complex mass order and the number of zero-energy bound states goes by the name of the Jackiw-Rossi index theorem [10, 11].
@@
The above-mentioned three index theorems, although originated in the context of high-energy physics, found their direct and important implications in various quantum materials, featuring emergent Dirac quasiparticles (massless or massive) as low-energy excitations around a few isolated points in the Brillouin zone, known as the Dirac or Weyl materials [12, 13, 14]. For example, the Jackiw-Rebbi index theorem directly applies to the Su-Schrieffer–Heeger chain in one dimension, when the effective Dirac Hamiltonian features a domain-wall mass [15, 16]. The same mechanism is also responsible for the existence of zero-energy one-dimensional edge modes and two-dimensional surface states at the boundaries of two- and three-dimensional strong topological insulators, respectively [17, 18]. The Aharonov-Casher index theorem, on the other hand, dictates the robustness of the zeroth Landau level for graphene, subject to external magnetic fields of arbitrary spatial profile [19, 20], which is applicable to Dirac fermions subject to chiral or axial magnetic field [21, 22, 23] as well as the ones residing on a negatively curved hyperbolic space [24]. Finally, the signature of the Jackiw-Rossi index theorem gets imprinted on the zero-energy modes near the vortex core of an $s$-wave superconductor otherwise realized on the surface of a three-dimensional topological insulator via proximity effect [25]. The same index theorem also applies to graphene when the ground state harbors a Kekulé valence-bond order [26], assisted by electron-phonon interaction [27] or an easy-plane antiferromagnetic order, assisted by on-site Hubbard repulsion and in-plane Zeeman coupling [28]. In both cases, the mass orders possess the requisite U(1) symmetry in the ordered states and zero-energy modes are found in the spectrum. Existence of such zero-energy modes leads to fascinating phenomena such as charge fractionalization [29, 30, 31, 32] and competing orders near the vortex core [33, 34].
@@
In this work, I set out to establish the jurisdiction of these three key index theorems to a class of quasi-relativistic open quantum systems, effectively described by non-Hermitian (NH) Dirac operators. In this pursuit, I strictly focus on a specific family of NH Dirac operators that enjoy all the quintessential elegance of the original Dirac theory [1, 2, 3]. Namely, such NH Dirac operators support guaranteed real eigenvalues over an extended NH parameter regime and display quasi-relativistic Lorentz invariance. Therefore, an introduction to the construction of such a NH Dirac theory would greatly facilitate the forthcoming summary of the central outcomes of this work, which I present in Sec. I.2.
@@ I.1 NH Dirac theory: Construction
The Dirac theory permits another class of Lorentz scalars, namely the Dirac masses, represented by a set of Hermitian matrices $\{M\}$ that fully anticommutes with $H_{\rm Dir}(\bm{k})$ and squares to the identity matrix. In terms of these two Lorentz invariant quantities, I define an anti-Hermitian operator $MH_{\rm Dir}(\bm{k})$, which also transforms as a scalar under the Lorentz transformation. The Lorentz invariant NH Dirac operator is then given by [35, 36, 37, 38, 39]
@@
In this work, I scrutinize the Aharonov-Casher, Jackiw-Rebbi, and Jackiw-Rossi index theorems within the framework of such NH Dirac operators. An important comment is due at this stage. In the last two cases, the Hermitian matrices describing the mass orders that assume spatially nontrivial topological textures (domain-wall and vortices) commute with $M$ from Eq. (2), such that they fully anticommute with $H^{\rm NH}_{\rm Dir}(\bm{k})$ and hence represent genuine mass orders for NH Dirac fermions. Only then the eigenspectrum of the corresponding total NH massive Dirac operator can be purely real over an extended NH parameter regime. Such mass orderings in the context of NH Dirac theory is named ‘commuting class mass’ [35], which will be discussed in Secs. III and IV. Next, I present a synopsis of the main results.
@@ I.2 Summary of main results
I begin the discussion by scrutinizing the fate of the Aharonov-Casher index theorem for planar massless NH Dirac fermions subject to external magnetic fields of arbitrary shape such that the system encloses $N$ number of magnetic flux quanta. Such a system is shown to host exactly $N$ number of normalizable zero-energy modes irrespective of the value of the NH parameter ($\alpha$) as their localization length is solely determined by the magnetic field profile and is completely insensitive to the effective Fermi velocity of NH Dirac fermions, see Sec. II.
@@
Next I establish the Jackiw-Rebbi index theorem for one-dimensional massive NH Dirac system in which the mass operator that assumes a domain-wall texture commutes with the mass matrix $M$ entering the construction of the NH Dirac operator in Eq. (2). Such mass orders are named ‘commuting class masses’. In such a system, I find localized and normalizable zero-energy modes at the core of the domain-wall only when the effective Fermi velocity of the collection of NH Dirac fermions is real. This outcome can be justified in the following way. The ratio of the magnitude of the asymptotic value of the mass as $x\to\pm\infty$ to the effective Fermi velocity of NH Dirac fermions determines the localization length of the zero-energy modes bound to the domain-wall defect core. Naturally, only when the effective Fermi velocity is real the localization length of zero modes is also real and they can be found in the spectrum of massive NH one-dimensional Dirac fermions. See Sec. III for details.
@@
Finally, I extend the jurisdiction of the Jackiw-Rossi index theorem for two-dimensional massive Dirac fermions in the presence of a composite mass order with the U(1) symmetry that assumes a vortex-type real space topological defect in a NH setup. The mass matrix $M$ in the construction of NH Dirac operator [Eq. (2)] is chosen such that it commutes with the composite mass order, which thus once again belongs to the family of commuting class masses. In this setup, I show that a mass vortex of integer vorticity $n$ supports exactly $n$ number of zero-energy modes only when the effective Fermi velocity of NH Dirac fermions is real. By exploiting a pseudo particle-hole symmetry of the NH Jackiw-Rossi operator, I extend this model to include additional terms therein yielding the generalized NH Jackiw-Rossi operator. In this construction, I show that the system supports only one zero-energy mode when $n$ is an odd integer, while the spectrum is devoid of any such zero modes for any even $n$. Thus, in terms of the number of zero-energy modes, the generalized NH Jackiw-Rossi operator features a $Z_{2}$ index protected by a pseudo particle-hole symmetry, while the conventional NH Jackiw-Rossi operator renders a $Z$ index, protected by a unitary particle-hole symmetry. See Sec. IV for details. I also note that when the Fermi velocity of NH Dirac fermions become purely imaginary, these zero-energy bound states tied to domain-wall and vortices become plane-waves at zero-energy, a phenomenon only occurs in NH Dirac systems.
@@
Although in this work I focus on three prominent index theorems for quasi-relativistic Dirac fermions in open or NH setups, based on the outcomes one can conjecture the following generic statement. The spectrum of Lorentz invariant NH Dirac fermions should support the same number of zero-energy modes as in Hermitian systems for arbitrary strength of the NH parameter if the localization length of such mid-gap bound states is independent of the effective Fermi velocity, as is the case for the Aharonov-Casher index theorem for gapless or massless Dirac fermions. On the other hand, if such bound states arise in the spectrum of massive Dirac fermions, then their number remains the same as in the Hermitian system, but only if the effective Fermi velocity is real, as is the case for Jackiw-Rebbi and Jackiw-Rossi index theorems. Finally, it should be noted that for massive Dirac fermions such index theorems hold in NH systems when the mass orders, assuming topologically nontrivial spatial textures (such as domain-wall and vortices, for example), belong to the ‘commuting class mass’ family.
@@ I.3 Organization
The remainder of the discussion is organized as follows. In Sec. II, I show the Aharonov-Casher index theorem for planar NH gapless Dirac fermions. Section III is devoted to demonstrate the Jackiw-Rebbi index theorem for NH one-dimensional massive Dirac fermions in the presence of a domain-wall of a commuting class mass. The Jackiw-Rossi index theorem of a composite mass order for two-dimensional NH Dirac systems is discussed in Sec. IV. A summary of results and possible experimental platform to test the theoretical predictions are staged in Sec. V.
@@ II Aharonov-Casher index theorem
The Aharonov-Casher index theorem concerns the number of zero-energy bound states when planar massless Dirac fermions experience uniform or non-uniform magnetic fields, such that the system encloses $N$ number of magnetic flux quanta. In this section, I establish this index theorem for a collection of NH massless Dirac fermions, described by the NH operator from Eq. (2). For concreteness, I consider its minimal two-component representation in $d=2$, for which $\Gamma_{1}=\sigma_{1}$ and $\Gamma_{2}=\sigma_{2}$. Then a natural and unique choice of $M$ is $M=\sigma_{3}$. Here, $\{\sigma_{\mu}\}$ is the set of two-dimensional Pauli matrices with $\mu=1,2,3$. The explicit form of the NH operator in the presence of external magnetic field reads as
@@
Here I show the explicit solutions for the right-eigenvectors $\left|\Psi_{0}\right>_{R}$ associated with the zero-energy modes, satisfying the secular equation
@@
where $\varphi$ is the polar angle, $C$ is the normalization constant, about which in a moment, and $j$ is an integer required for the single-valued right wavefunctions. The spatial profile of the external magnetic field $B(\bm{r})$ is obtained from the Green’s function associated with the Laplacian in two dimensions, leading to
@@
where $\Phi=\int d^{2}\bm{r}^{\prime}B(\bm{r}^{\prime})$ is the total magnetic flux enclosed by the system. Also note that in the natural units ($\hbar=1$) the magnetic flux quantum $\Phi_{0}=h/e\to 2\pi/e$ and thus $e\Phi/2\pi\equiv\Phi/\Phi_{0}=N$ is the total magnetic flux quanta enclosed by the system.
@@
Once again the NH parameter ($\alpha$) does not enter the solution for the left-eigenvectors for the zero modes, for which I find
@@
The normalization condition is then fixed by the bi-orthonormal condition, given by [40]
@@
for which the solutions remain normalizable. Notice that the $N$ allowed values of $j$ are completely independent of the profile of the magnetic field, and are solely determined by the total magnetic flux quanta $N$ enclosed by the planar NH quasi-relativistic system. Furthermore, these solutions can always be found for any arbitrary value of $\alpha$ or the effective Fermi velocity $v_{{}_{\rm F}}$ of the collection of NH Dirac fermions. This is so because the localization length of the zero modes is solely determined by the magnetic flux profile in which $v_{{}_{\rm F}}$ plays no role, see Eqs. (9) and (13). Hence, the Aharonov-Casher index theorem apply to quasi-relativistic massless NH Dirac fermions in two dimensions, irrespective of the strength of non-Hermiticity in the system. Next, I discuss two cases where the reality condition of the effective Fermi velocity of NH Dirac fermions plays a decisive role for the existence of normalizable zero-energy modes.

### v2

@@ Abstract
Dirac fermions, subject to external magnetic fields and in the presence of mass orders that assume topologically nontrivial spatial textures such as domain wall and vortices, for example, bind robust midgap states at zero energy, the number of which is governed by the Aharonov-Casher and Jackiw-Rebbi or Jackiw-Rossi index theorems, respectively. Here I extend the jurisdiction of these prominent index theorems to Lorentz invariant non-Hermitian (NH) Dirac operators, constructed by augmenting the celebrated Dirac Hamiltonian by a masslike anti-Hermitian operator that also scales linearly with momentum. The resulting NH Dirac operator manifests all-real eigenvalues over an extended NH parameter regime, characterized by a real effective Fermi velocity for NH Dirac fermions ($v_{{}_{\rm F}}$). From the explicit solutions of the zero-energy bound states, I show that in the presence of external magnetic fields of arbitrary shape such modes always exist when the system encloses a finite number of magnetic flux quanta, while in the presence of spatially nontrivial textures of the mass orders localized zero-energy modes can only be found in the spectrum when $v_{{}_{\rm F}}$ is real. These findings pave a concrete route to realize nucleation of competing orders from the topologically robust zero-energy manifold in NH or open Dirac systems. Possible extensions of these outcomes to other index theorems and tabletop experimental setups to test these predictions are discussed.
@@ I Introduction
Zero energy in the Dirac theory enjoys a special status as the associated eigenvalue spectrum extends equally to positive and negative values [1, 2, 3]. Therefore, any bound state that resides precisely at zero energy often (if not always) is robust due to the intrinsic particle-hole symmetry in the system. Such a situation occurs at least in two specific cases. (a) When massless Dirac fermions, confined to an Euclidean plane, are subject to external uniform or inhomogeneous magnetic fields, the number of zero-energy bound states is exactly equal to the total magnetic flux quanta enclosed by the system, guaranteed by the Aharonov-Casher index theorem [4, 5, 6]. (b) Bound states at zero energy appear in the spectrum of Dirac fermions when they foster mass orders, captured by Hermitian operators fully anticommuting with the Dirac Hamiltonian, that assume spatially nontrivial topological textures. For example, in one spatial dimension, when the Dirac mass takes the shape of a domain wall a zero-energy mode gets pinned where the mass order changes sign, a result known as the Jackiw-Rebbi index theorem [7, 8, 9]. In two spatial dimensions, when a composite Dirac mass, described by two mutually anticommuting mass matrices with the requisite U(1) symmetry between them, takes the texture of a vortex, $n$ number of zero-energy bound states get pinned near the core of such topological defects, where $n$ is the vorticity of the mass texture. Such a one-to-one correspondence between the vorticity of the complex mass order and the number of zero-energy bound states goes by the name of the Jackiw-Rossi index theorem [10, 11].
@@
The above-mentioned three index theorems, although originated in the context of high-energy physics, found their direct and important implications in various quantum materials, featuring emergent Dirac quasiparticles (massless or massive) as low-energy excitations around a few isolated points in the Brillouin zone, known as the Dirac or Weyl materials [12, 13, 14]. For example, the Jackiw-Rebbi index theorem directly applies to the Su-Schrieffer–Heeger chain in one dimension, when the effective Dirac Hamiltonian features a domain-wall mass [15, 16]. The same mechanism is also responsible for the existence of zero-energy one-dimensional edge modes and two-dimensional surface states at the boundaries of two- and three-dimensional strong topological insulators, respectively [17, 18], and zero-energy flat band in uniaxially strained honeycomb lattice [19]. The Aharonov-Casher index theorem, on the other hand, dictates the robustness of the zeroth Landau level for graphene, subject to external magnetic fields of arbitrary spatial profile [20, 21], which is applicable to Dirac fermions subject to chiral or axial magnetic fields [22, 23, 24] as well as the ones residing on a negatively curved hyperbolic space [25]. Finally, the signature of the Jackiw-Rossi index theorem gets imprinted on the zero-energy modes near the vortex core of an $s$-wave superconductor otherwise realized on the surface of a three-dimensional topological insulator via proximity effect [26]. The same index theorem also applies to graphene when the ground state harbors a Kekulé valence-bond order [27], assisted by electron-phonon interaction [28] or an easy-plane antiferromagnetic order, assisted by on-site Hubbard repulsion and in-plane Zeeman coupling [29]. In both cases, the mass orders possess the requisite U(1) symmetry in the ordered states and zero-energy modes are found in the spectrum. Existence of such zero-energy modes leads to fascinating phenomena such as charge fractionalization [30, 31, 32, 33] and competing orders near the vortex core [34, 35, 36].
@@
In this work, I set out to establish the footprints of these three key index theorems on a class of quasirelativistic open quantum systems, effectively described by non-Hermitian (NH) Dirac operators. In this pursuit, I strictly focus on a specific family of NH Dirac operators that enjoy all the quintessential elegance of the original Dirac theory [1, 2, 3]. Namely, such NH Dirac operators support guaranteed all-real eigenvalues over an extended NH parameter regime and display quasirelativistic Lorentz invariance. Therefore, an introduction to the construction of such a NH Dirac theory would greatly facilitate the forthcoming summary of the central outcomes of this work, which I present in Sec. I.2.
@@ I.1 NH Dirac theory: Construction
The Dirac theory permits another class of Lorentz scalars, namely the Dirac masses, represented by a set of Hermitian matrices $\{M\}$ that fully anticommutes with $H_{\rm Dir}(\bm{k})$ and squares to the identity matrix [37]. In terms of these two Lorentz invariant quantities, I define an anti-Hermitian operator $MH_{\rm Dir}(\bm{k})$, which also transforms as a scalar under the Lorentz transformation. The Lorentz invariant NH Dirac operator is then given by [38, 39, 40, 41, 42]
@@
In this work, I scrutinize the Aharonov-Casher, Jackiw-Rebbi, and Jackiw-Rossi index theorems within the framework of such NH Dirac operators. An important comment is due at this stage. In the last two cases, the Hermitian matrices describing the mass orders that assume spatially nontrivial topological textures (domain wall and vortices) are described by Hermitian operators that commute with $M$ from Eq. (2) and anticommute with $H_{\rm Dir}(\bm{k})$ from Eq. (1), such that they fully anticommute with anti-Hermitian operator $MH_{\rm Dir}(\bm{k})$ and thus anticommute with the total NH operator $H^{\rm NH}_{\rm Dir}(\bm{k})$, hence representing genuine mass orders for NH Dirac fermions. Only then the eigenspectrum of the corresponding total NH massive Dirac operator can be purely real over an extended NH parameter regime, as then the effective single-particle Hamiltonian in the ordered phase is captured by matrix operators all of which mutually anticommute with each other. Such mass ordering in the context of NH Dirac theory is named “commuting class masses” [38], which will be discussed in Secs. III and IV. Next, I present a synopsis of the main results.
@@ I.2 Summary of main results
I begin the discussion by scrutinizing the fate of the Aharonov-Casher index theorem for planar massless NH Dirac fermions subject to external perpendicular magnetic fields of arbitrary shape such that the system encloses $N$ number of magnetic flux quanta. Such a system is shown to host exactly $N$ number of normalizable zero-energy modes irrespective of the value of the NH parameter ($\alpha$) as their localization length is solely determined by the magnetic field profile and is completely insensitive to the effective Fermi velocity of NH Dirac fermions ($v_{{}_{\rm F}}$). See Sec. II for details.
@@
Next, I establish the Jackiw-Rebbi index theorem for a one-dimensional massive NH Dirac system in which the mass operator that assumes a domain-wall texture commutes with the mass matrix $M$ entering the construction of the NH Dirac operator in Eq. (2). Such mass orders are named “commuting class masses.” In such a system, I find localized and normalizable zero-energy modes at the core of the domain wall only when the effective Fermi velocity of the collection of NH Dirac fermions is real. This outcome can be justified in the following way. The ratio of the magnitude of the asymptotic value of the mass as $x\to\pm\infty$ to the effective Fermi velocity of NH Dirac fermions determines the localization length of the zero-energy modes bound to the domain-wall defect core. Naturally, only when the effective Fermi velocity is real the localization length of zero modes is also real and they can be found in the spectrum of massive NH one-dimensional Dirac fermions. See Sec. III for details.
@@
Finally, I extend the jurisdiction of the Jackiw-Rossi index theorem for two-dimensional massive Dirac fermions in the presence of a composite mass order with the U(1) symmetry that assumes a vortex-type real space topological defect in a NH setup. The mass matrix $M$ in the construction of the NH Dirac operator [Eq. (2)] is chosen such that it commutes with the composite mass order, which thus once again belongs to the family of “commuting class masses.” In this setup, I show that a mass vortex of integer vorticity $n$ supports exactly $n$ number of zero-energy modes only when the effective Fermi velocity of NH Dirac fermions is real. By exploiting a pseudo-particle-hole symmetry of the NH Jackiw-Rossi operator, I extend this model to include additional terms therein yielding the generalized NH Jackiw-Rossi operator, that, for example include the orbital coupling of a gauge field required for the stability of isolated or deconfined vortices. In this construction, I show that the system supports only one zero-energy mode when $n$ is an odd integer, while the spectrum is devoid of any such zero modes for any even $n$. Thus, in terms of the number of zero-energy modes, the generalized NH Jackiw-Rossi operator features a $Z_{2}$ index protected by a pseudo-particle-hole symmetry, while the conventional NH Jackiw-Rossi operator renders a $Z$ index, protected by a unitary particle-hole symmetry. See Sec. IV for details. I also note that when the Fermi velocity of NH Dirac fermions becomes purely imaginary, i.e. when $|\alpha|>1$, these zero-energy bound states tied to domain-wall and vortices become plane waves at zero-energy, a phenomenon that only occurs in NH Dirac systems.
@@
Although in this work I focus on three prominent index theorems for quasirelativistic Dirac fermions in open or NH setups, based on the outcomes one can conjecture the following generic statement. The spectrum of Lorentz invariant NH Dirac fermions should support the same number of zero-energy modes as in Hermitian systems for arbitrary strength of the NH parameter if the localization length of such midgap bound states is independent of the effective Fermi velocity, as is the case for the Aharonov-Casher index theorem for gapless or massless Dirac fermions. On the other hand, if such bound states arise in the spectrum of massive Dirac fermions, then their number remains the same as in the Hermitian system, but only if the effective Fermi velocity is real, as is the case for Jackiw-Rebbi and Jackiw-Rossi index theorems. Finally, it should be noted that for massive Dirac fermions such index theorems hold in NH systems when the mass orders, assuming topologically nontrivial spatial textures (such as domain wall and vortices, for example), belong to the “commuting class mass” family.
@@ I.3 Organization
The remainder of the discussion is organized as follows. In Sec. II, I show the Aharonov-Casher index theorem for planar NH gapless Dirac fermions, subject to external perpendicular magnetic fields. Section III is devoted to demonstrating the Jackiw-Rebbi index theorem for NH one-dimensional massive Dirac fermions in the presence of a domain-wall of a “commuting class mass.” The Jackiw-Rossi index theorem of a composite mass order for two-dimensional NH Dirac systems is discussed in Sec. IV. A summary of results, their possible extensions, and promising tabletop experimental platforms to test the theoretical predictions are staged in Sec. V.
@@ II Aharonov-Casher index theorem
The Aharonov-Casher index theorem concerns the number of zero-energy bound states when planar massless Dirac fermions experience uniform or nonuniform perpendicular magnetic fields, such that the system encloses $N$ number of magnetic flux quanta. In this Section, I establish this index theorem for a collection of NH massless Dirac fermions, described by the NH operator from Eq. (2). For concreteness, I consider its minimal two-component representation in $d=2$, for which $\Gamma_{1}=\sigma_{1}$ and $\Gamma_{2}=\sigma_{2}$. Then a natural and unique choice of $M$ is $M=\sigma_{3}$. Here, $\{\sigma_{\mu}\}$ is the set of two-dimensional Pauli matrices with $\mu=1,2,3$. The explicit form of the NH operator in the presence of an external magnetic field reads
@@
Here I show the explicit solutions for the right eigenvectors $\left|\Psi_{0}\right>_{R}$ associated with the zero-energy modes, satisfying the secular equation
@@
where $\varphi$ is the polar angle, $C$ is the normalization constant, discussed in a moment, and $j$ is an integer required for the single-valued right wave functions. The scalar function $\chi(\bm{r})$ is obtained from the spatial profile of the external magnetic field $B(\bm{r})$ via the Green’s function associated with the Laplacian in two dimensions as
@@
where $\Phi=\int d^{2}\bm{r}^{\prime}B(\bm{r}^{\prime})$ is the total magnetic flux enclosed by the system. Notice that in the natural units ($\hbar=1$) the magnetic flux quantum $\Phi_{0}=h/e\to 2\pi/e$ and thus $e\Phi/2\pi\equiv\Phi/\Phi_{0}=N$ is the total magnetic flux quanta enclosed by the system.
@@
Once again the NH parameter ($\alpha$) does not enter the solution for the left eigenvectors for the zero modes, for which I find
@@
The normalization condition is then fixed by the biorthonormal condition, given by [43]
@@
for which the solutions remain normalizable. Notice that the $N$ allowed values of $j$ are completely independent of the profile of the magnetic field, and are solely determined by the total magnetic flux quanta $N$ enclosed by the planar NH quasirelativistic system. Furthermore, these solutions can always be found for any arbitrary value of $\alpha$ or the effective Fermi velocity $v_{{}_{\rm F}}$ of the collection of NH Dirac fermions. This is so because the localization length of the zero modes is solely determined by the magnetic flux profile in which $v_{{}_{\rm F}}$ plays no role, see Eqs. (9) and (13). Hence, the Aharonov-Casher index theorem applies to quasirelativistic massless NH Dirac fermions in two dimensions, irrespective of the strength of non-Hermiticity in the system. Next, I discuss two cases where the reality condition of the effective Fermi velocity of NH Dirac fermions plays a decisive role for the existence of normalizable zero-energy modes.